# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

### v2

@@ Abstract
Inferring trajectories from longitudinal spatially-resolved omics data is fundamental to understanding the dynamics of structural and functional tissue changes in development, regeneration and repair, disease progression, and response to treatment. We propose ContextFlow, a novel context-aware flow matching framework that incorporates prior knowledge to guide the inference of structural tissue dynamics from spatially resolved omics data. Specifically, ContextFlow integrates local tissue organization and ligand-receptor communication patterns into a transition plausibility matrix that regularizes the optimal transport objective. By embedding these contextual constraints, ContextFlow generates trajectories that are not only statistically consistent but also biologically meaningful, making it a generalizable framework for modeling spatiotemporal dynamics from longitudinal, spatially resolved omics data. Comprehensive experiments show that ContextFlow consistently outperforms state-of-the-art flow matching methods across multiple metrics of inference accuracy and biological coherence. Our code is available at: https://github.com/santanurathod/ContextFlow.
Machine Learning, ICML
@@ 1 Introduction
Flow matching (Lipman et al., 2023) is an emerging generative paradigm that provides an efficient approach for learning the complex latent dynamics, or normalizing flows (Papamakarios et al., 2021), of a system of variables, while enabling parametric flexibility to model data distributions. Inferring the underlying dynamics from sparse and noisy observations is a central challenge in many application domains (Gontis et al., 2010; Brunton et al., 2016; Pandarinath et al., 2018; Li et al., 2025), where continuous trajectories are rarely captured; instead, cross-sectional snapshots, collected at discrete time points, are typically available. In single-cell RNA sequencing (scRNA-seq), this challenge becomes especially critical as the destructive nature of profiling technologies yields only unpaired population-level snapshots over time. Uncovering temporal dynamics from such snapshot data is essential for understanding developmental processes, disease progression, treatment, and perturbation responses (Wagner and Klein, 2020). Traditional approaches often rely on heuristics or computationally intensive likelihood-based generative models, which struggle with scalability and flexibility in high-dimensional single-cell data. In contrast, flow matching overcomes these challenges by directly learning continuous latent dynamics that are constrained to match observed population-level distributions at sampled time points.
Table 1: Comparisons of state-of-the-art spatiotemporal OT and flow matching methods. In particular, ContextFlow excels at efficiently inferring biologically plausible trajectories by incorporating spatial priors into minibatch-OT-based flow-matching frameworks.
Method Generative Prior Knowledge Unified Velocity Field OT Runtime
DeST-OT (Halmos et al., 2025) $O(N^{3})$
TOAST (Ceccarelli et al., 2025) $O(N^{3})$
PASTE (Zeira et al., 2022) $O(N^{3})$
CFM (Lipman et al., 2022)
MOTFM (Tong et al., 2024) $O(N^{2})$
ContextFlow (ours) $O(N^{2})$
@@
The state and function of cells within a tissue are affected by interactions with neighboring cells, extracellular matrix components, and local signaling gradients (Rao et al., 2021). Recent advances in spatial omics technologies, particularly spatial transcriptomics (ST), allow gene expression profiling without tissue dissociation, thereby preserving spatial context and providing a complementary view of cellular organization. The dynamics of complex cellular processes are affected by the tissue microenvironment, where cells engage in reciprocal communication with their neighbors (Dimitrov et al., 2022; Tanevski et al., 2025). A growing body of work highlights the critical role of spatial cell–cell communication patterns in shaping cellular phenotypes (Armingol et al., 2021). In particular, location-specific communication circuits between distinct cell types dynamically interact to reprogram cellular states and influence tissue-level behavior (Mayer et al., 2023; Aguadé-Gorgorió et al., 2024; Zheng et al., 2025). These insights, made possible by the spatiotemporal resolution of transcriptomics data, pave the way for understanding the mechanisms by which cellular interactions drive tissue organization and function in organogenesis (Chen et al., 2022), regeneration (Ben-Moshe et al., 2021; Wei et al., 2022), disease progression (Kukanja et al., 2024), and treatment response (Liu et al., 2024).
Optimal transport (OT) has become a foundational framework for aligning spatially resolved samples and inferring putative developmental or temporal couplings  (Zeira et al., 2022; Liu et al., 2023). As a result, state-of-the-art flow matching frameworks such as minibatch-OT flow matching (MOTFM) (Tong et al., 2024) use OT-derived couplings to define conditional paths for training velocity fields, thereby overcoming the lack of generative capabilities in optimal transport. The OT formulation adopted in MOTFM, however, does not account for the contextual richness of spatial transcriptomics and can yield trajectories that are statistically optimal yet biologically implausible (see Figure 4(a) in Appendix G.1 for an illustration). While recent studies (Halmos et al., 2025; Ceccarelli et al., 2025) have extended OT objectives to spatial transcriptomics, they primarily focus on pairwise alignment of populations across conditions or modalities and do not explicitly incorporate the cell–cell communication patterns that drive cellular state transitions.

### v3

@@ Abstract
Inferring trajectories from longitudinal, spatially resolved omics data is fundamental to understanding the dynamics of structural and functional tissue changes in development, regeneration and repair, disease progression, and response to treatment. We propose ContextFlow, a context-aware flow matching framework that leverages prior knowledge to guide the inference of structural tissue dynamics from spatially resolved omics data. Existing approaches rely on optimal-transport-derived couplings that are determined solely by transcriptomic similarity, which can lead to biologically implausible transitions and, in turn, inconsistent learned dynamics. Specifically, ContextFlow integrates local tissue organization and ligand-receptor communication patterns into a transition plausibility matrix that regularizes the optimal transport objective. By embedding these contextual constraints, ContextFlow generates trajectories that are not only statistically consistent but also biologically meaningful, making it a generalizable framework for modeling cellular dynamics from longitudinal, spatially resolved omics data. Comprehensive experiments show that ContextFlow consistently improves state-of-the-art flow matching methods across multiple metrics of inference accuracy and biological coherence. Our code is available at: https://github.com/santanurathod/ContextFlow.
@@ 1 Introduction
Inferring the underlying dynamics from sparse and noisy observations is a central challenge in many applications (Gontis et al., 2010; Brunton et al., 2016; Pandarinath et al., 2018; Li et al., 2025), where continuous trajectories are rarely captured; instead, cross-sectional snapshots, collected at discrete time points, are available. In single-cell RNA sequencing, this challenge becomes especially critical, as the destructive nature of profiling technologies yields only unpaired population-level snapshots over time. Uncovering temporal dynamics from such snapshot data is essential for understanding developmental processes, disease progression, treatment, and perturbation responses (Wagner and Klein, 2020). Traditional methods (Trapnell et al., 2017; Chen et al., 2018; Rubanova et al., 2019; Bergen et al., 2020) rely on heuristics or computationally intensive likelihood-based generative models, which struggle with flexibility and scalability in handling high-dimensional data. In contrast, flow matching (Lipman et al., 2023), an emerging generative paradigm, overcomes these issues by directly learning continuous latent dynamics constrained to match population-level distributions at observational time points, while enabling parametric flexibility to model data distributions.
Cellular dynamics are affected by the tissue microenvironment, in which cells engage in reciprocal communication with their neighbors (Dimitrov et al., 2022; Tanevski et al., 2025). State-of-the-art flow matching frameworks (Tong et al., 2024; Kapuśniak et al., 2024; Atanackovic et al., 2024; Rohbeck et al., 2025) use optimal transport (OT) to derive couplings between consecutive observational time points and design conditional paths for regression over the velocity fields. Although OT couplings improve training stability and inference efficiency over vanilla independent couplings (Lipman et al., 2023), they are determined solely by transcriptomic similarity and do not account for the contextual richness of spatial transcriptomics (Armingol et al., 2021; Mayer et al., 2023; Aguadé-Gorgorió et al., 2024; Zhang et al., 2025b), thus yielding couplings that are statistically optimal but biologically implausible. Since OT-based flow matching algorithms directly regress on these couplings, these errors propagate into the learned velocity field, leading to globally inconsistent inferred dynamics.
To address these limitations, we introduce ContextFlow, a generative framework that incorporates explicit spatial priors into OT-coupled flow matching to model temporal tissue dynamics (see Figure 2 for an illustration). In contrast to existing approaches, ContextFlow integrates prior knowledge directly into the construction of OT couplings to avoid biologically implausible transitions as the regression target. By encoding local tissue organization and ligand–receptor-derived spatial communication patterns into prior-regularized optimal transport formulations, ContextFlow exploits the contextual richness of spatial omics data and embeds both structural and functional aspects of tissue organization into the training objective, thereby generating more biologically informed trajectories.
In summary, our contributions are as follows:
  • We define transitional plausibility to capture local tissue organization and ligand–receptor communication patterns, which provides a flexible and interpretable mechanism for encoding spatial and functional priors into couplings used to train the velocity field (Section 4.1). Our definition allows priors to be explicitly and flexibly adapted in a domain-specific manner, enabling the incorporation of external biological knowledge and improving interpretability.
  • We introduce principled methods for integrating explicit biological priors into the state-of-the-art OT-based flow matching framework, instantiated via two novel schemes—cost-based and entropy-based—both amenable to efficient Sinkhorn optimization (Section 4.2).
  • Comprehensive experiments on regeneration and developmental datasets show that ContextFlow consistently improves existing methods for both interpolation and extrapolation tasks across metrics that capture biological plausibility and statistical fidelity (Section 5).
2 Related Work
2.1 Flow Matching and its Applications in Biology
Flow matching (FM) (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2022; Liu et al., 2022) is a generative modeling framework that trains the velocity field via a tractable regression objective for learning continuous normalizing flows (Papamakarios et al., 2021), which is more efficient and scalable when dealing with high-dimensional data than prior approaches, such as neural ODE (Chen et al., 2018; Rubanova et al., 2019). Compared with score-based diffusion models (Song et al., 2020) and Schrödinger Bridges (SB) (Bunne et al., 2023), FM is faster and more stable to train and sample from during inference, as it characterizes the latent continuous dynamics via an ODE. Therefore, a growing body of literature has adapted and improved the vanilla FM framework to model dynamic scientific processes, among others, especially in biology and the life sciences (Li et al., 2025).
Instead of independent couplings, (Pooladian et al., 2023; Tong et al., 2024) used optimal transport (OT) maps to define the conditional coupling between the source and the target, and approximated it using minibatch samples, leading to more stable training and faster inference via straightening the flows. Based on this framework,  (Kapuśniak et al., 2024) introduced metric flow matching to learn the flows over a non-Euclidean manifold to make a case against straighter flows in the context of transcriptomics. (Atanackovic et al., 2024) proposed meta flow matching, which incorporates implicit cell-cell interaction prior via GNN conditioning. Apart from these trajectory inference-centric studies, FM has also been utilized to generate imaging-based cell morphology changes (Zhang et al., 2025a), simulate spatial transcriptomics data from histology images (Huang et al., 2025), and approximate OT maps for drug response modeling and cross-modal translation tasks Klein et al. (2024). Despite these advances, existing work does not address how to meaningfully incorporate explicit biological prior knowledge to constrain the velocity field, thereby limiting the biological plausibility of inferred trajectories, interpretability, and control over the learned dynamics.
2.2 Advancement in Spatial Transcriptomics
@@
The state and function of cells within a tissue are influenced by interactions with neighboring cells, extracellular matrix components, and local signaling gradients (Rao et al., 2021). Recent advances in spatial omics technologies, particularly spatial transcriptomics (ST), allow gene expression profiling without tissue dissociation, thereby preserving spatial context and providing a complementary view of cellular organization. The dynamics of cellular processes are regulated by the tissue microenvironment and by cell communications with their neighbours (Dimitrov et al., 2022; Tanevski et al., 2025). A growing body of literature highlights the critical role of spatial cell–cell communication patterns in shaping cellular phenotypes (Armingol et al., 2021). In particular, location-specific communication circuits between distinct cell types dynamically interact to reprogram cellular states and influence tissue-level behavior (Mayer et al., 2023; Aguadé-Gorgorió et al., 2024; Zheng et al., 2025). These insights, made possible by the spatiotemporal resolution of transcriptomics data, pave the way for understanding the mechanisms by which cellular interactions drive tissue organization and function in organogenesis (Chen et al., 2022), regeneration (Ben-Moshe et al., 2021; Wei et al., 2022), disease progression (Kukanja et al., 2024), and treatment response (Liu et al., 2024).
Optimal transport (Bunne et al., 2024; Klein et al., 2025) has been increasingly used to analyze omics data, as they often exhibit uncoupled measurements across modalities, conditions, or time points. In ST, several OT formulations have been introduced depending on context. For instance, Zeira et al. (2022) and Liu et al. (2023) proposed PASTE and PASTE2 to align ST data from adjacent tissue slices, while DeST-OT (Halmos et al., 2025) integrates spatio-temporal slices by modeling cell growth and differentiation. In addition, Rahimi et al. (2024) developed DOT, a multi-objective OT framework for mapping features across scRNA-seq and spatially resolved assays, and Ceccarelli et al. (2026) introduced TOAST, a spatially regularized OT framework for slice alignment and annotation transfer. While these OT techniques can successfully leverage spatial or structural information to improve alignment quality across space, time, and modalities, they do not learn a velocity field governing temporal evolution, as in generative modeling frameworks like flow matching. Consequently, while these techniques can model temporal relationships and align distributions across time, they do not learn a continuous-time generative model and therefore cannot directly simulate intermediate or future cellular states. Moreover, although some of the aforementioned approaches incorporate spatial or structural constraints, these are typically introduced in a task-specific or implicit manner, rather than as general mechanisms to encode explicit biological priors directly into the transport plan.
3 Problem Formulation
We aim to infer the temporal dynamics of cell states from spatially resolved gene expression data. Let $0=t_{1}<t_{2}<\ldots<t_{m+1}=1$ be a sequence of normalized time points and $q_{i}$ be the underlying distribution over $\mathbb{R}^{d}$ at time $t_{i}$. Given $\{\mathbf{X}_{t_{i}}\}_{i\in\{1,2,\ldots,m+1\}}$, where $\mathbf{X}_{t_{i}}=\{\bm{x}_{i}(k)\}_{k=1}^{n_{i}}$ corresponds to the gene expressions at time $t_{i}$ consisting of $n_{i}$ snapshot data sampled from $q_{i}$, our goal is to learn a neural velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to faithfully characterize the temporal evolution of spatially resolved tissues over time, such that the induced probability path $p_{t}$ can describe the state of each cell at any time $t\in[0,1]$. A standard approach is to apply conditional flow matching with minibatch OT couplings (MOTFM) (Tong et al., 2024) for each pair of consecutive time points. Specifically, for any $t\in[0,1]$ satisfying $t\in[t_{i},t_{i+1})$, the following linear conditional velocity field $u_{t}(\bm{x}|\bm{z})$ and its associated probability path $p_{t}(\bm{x}|\bm{z})$ are defined as the regression target in MOTFM:
$u_{t}(\bm{x}|\bm{z}_{i})=\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}}\quad\text{and}\quad p_{t}(\bm{x}|\bm{z}_{i})=\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right),$ (1)
where $\bm{z}_{i}=(\bm{x}_{i},\bm{x}_{i+1})$. Let $q(\bm{z}_{i})$ be the joint probability measure with marginals $q_{i}$ and $q_{i+1}$, corresponding to the entropic optimal transport (EOT) coupling, defined as:
$q(\bm{z}_{i})=\operatorname*{arg\,min}_{\pi\in\Pi(q_{i},q_{i+1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})+\epsilon H(\pi\>|\>q_{i}\otimes q_{i+1}),$ (2)
where $\epsilon>0$ is the regularization parameter, and $H(\pi\>|\>q_{i}\otimes q_{i+1})$ is the KL divergence with respect to the transport plan $\pi$ and the product measure $q_{i}\otimes q_{i+1}$. To train the velocity vector field $u_{\theta}$, MOTFM randomly samples a mini-batch of data at each time, uses the Sinkhorn algorithm (Cuturi, 2013) to obtain the EOT couplings, and optimizes the following regression loss using SGD:
$\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),\bm{z}\sim q(\bm{z}_{i}),\bm{x}\sim p_{t}(\bm{x}|\bm{z}_{i})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x}|\bm{z}_{i})\big\|^{2},$ (3)
where $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $i=i(t)$ denotes the integer that satisfies $t_{i}\leq t<t_{i+1}$ for each sampled $t$. Full preliminaries of flow matching are provided in Appendix B.
Despite their enhanced ability to model system dynamics, state-of-the-art OT-based flow matching frameworks lack provisions to fully exploit the contextual richness and integrate biological prior knowledge inferred from other associated data modalities. Existing approaches can generate statistically optimal trajectories by targeting probability paths induced by OT couplings along the temporal dimension. However, because they overlook important functional or structural prior information, their velocity fields can be trained on couplings that violate domain principles, leading to biologically implausible trajectories. Figure 1 presents an example of statistically optimal yet biologically implausible couplings between inhibitory-excitatory neuron cells coming from the EOT plan used to train MOTFM (Tong et al., 2024) (see Appendix G.1 for detailed discussions). This motivates our work to develop a mechanism to integrate domain prior knowledge into the OT-based flow matching framework, thereby encouraging domain-aligned training of the velocity field.
Figure 1: Visualization of inhibitory-exitatory coupled cells sampled from the EOT plan used in MOTFM on temporally consecutive spatial omics slides from the Brain Regeneration dataset (Wei et al., 2022). Here, different colors correspond to different cell types. Since inhibitory/excitatory cells serve distinct biological functions, their coupling and thus the implied state transition are infeasible.
4 Regularizing the Flow with Spatial Priors
4.1 Spatial Priors and Transitional Plausibility
To address the limitations, we propose modeling the context and cellular organization of spatial omics data using two types of spatial priors (spatial smoothness and cell-cell communication patterns), strong indicators of transitional plausibility between locations and cell states across time points.
Spatial Smoothness. Tissues are well-organized systems. Within a microenvironment, neighboring cells respond to the same set of external mechanical stimuli and intercellular communication, which affect their states similarly and result in local smoothness in cell-type-specific expression. Due to tissue heterogeneity, we cannot assume a common reference coordinate frame across tissue samples or even slices at $t_{i}$ and $t_{j}$ at a larger scale. However, the same heterogeneity allows us to consider spatial coherence and neighborhood consistency (Greenwald et al., 2024; Ceccarelli et al., 2026) as proxies for relative cell localization, which cannot change significantly over short time intervals. Therefore, the aggregate expression within each cell’s microenvironment can be used to quantify the plausibility of transition across consecutive time points.
$\mathrm{SS}(c_{i},c_{j})$ $=\bigg\|\frac{1}{|\mathcal{N}_{r}(c_{i})|}\sum_{c\in\mathcal{N}_{r}(c_{i})}\bm{x}(c)-\frac{1}{|\mathcal{N}_{r}(c_{j})|}\sum_{c\in\mathcal{N}_{r}(c_{j})}\bm{x}(c)\bigg\|_{2}^{2},$ (4)
where $\mathcal{N}_{r}(c_{i})=\{c:\|\bm{s}(c)-\bm{s}(c_{i})\|_{2}\leq r\}$ denotes the set of neighboring cells of $c_{i}$ in the same tissue slice, $|\mathcal{N}_{r}(c_{i})|$ is the cardinality of $\mathcal{N}_{r}(c_{i})$, and $\bm{x}(c)$ is the gene expression profile of cell $c$.
Cell-Cell Communication Patterns. Cell–cell communication plays a critical role in regulating biological processes, including development, apoptosis, and homeostasis in health and disease (Armingol et al., 2024). A major type of cell-cell communication is ligand–receptor (LR) signaling, in which ligands expressed by one cell bind to cognate receptors on another, initiating intracellular cascades that ultimately affect the expression profiles of the cell (Armingol et al., 2021). Numerous databases of prior knowledge of LR binding exist, and computational methods use them to systematically link gene expression with the activity of ligand-receptor-mediated communication.
Specifically, we represent each cell $c_{i}$ by a feature vector $f_{\mathrm{LR}}\in\mathbb{R}^{p}$, where each entry corresponds to one of $p$ possible ligand–receptor pairs and encodes the extent of $c_{i}$’s participation in communication through that pair. The transitional plausibility $\mathrm{TP}(c_{i},c_{j})$ between cells in different tissue slices is higher when they exhibit similar ligand-receptor communication patterns $f_{\mathrm{LR}}$ (see Figure 9 in Appendix G.3 for an illustration). Formally, we define $\mathrm{LR}(c_{i},c_{j})$, the dissimilarity between the ligand–receptor communication patterns in the microenvironments of cells $c_{i}$ and $c_{j}$, as: